Find the standard deviation of the following group of numbers (use n – 1 instead of n)…2, 5, 8, 9

6

3.14

2.74

3.16
Answer #1

Solution :

n = 4

\bar x = \sum x / n = (2 + 5 + 8 + 9) / 4 = 24 / 4 = 6

s = \sqrt \sum(x - \bar x)2 / n - 1  

= \sqrt [ (2 - 6)2 +(5 - 6) 2 + (8 - 6)2 + (9 - 6)2 ​​​​​​ / 4 - 1 ]

= ​​​​​​​3.16